Grade 12Mechanics

a block of mass 10 kg is moving in x direction with a constant speed of 10m/s. it is subjected to a retarding force F=-0.1x joule/metre during its travel from x=20 metre to x=30 metre. its final kinetic energy will be
  1. 475 J
  2. 450 J
  3. 275 J
  4. 250 J

silky srivastava
8 Years agoGrade 12

1 Answer

Apurva Sharma
ApprovedApproved Tutor Answer8 Years ago
The anser is (1)Final K.E= Iinitial KE -loss in KEInitial KE=1/2 × 10×10^2 =500 JNow, F at x=20 = 0.1×20= 2NF at x=30= 0/1×30= 3NAvg F= 2+3/5 =2.5NWork done=change in KE(loss)-> F.s= 2.5(30-20) =25JThus, Final KE = (500-25)J = 475J